# The mantle-inner core gravitational mode of oscillation in a strong magnetic field regime

### v1

@@ 2.2 The MICG mode
where $C_{tc}=C_{i}+C_{f}^{tc}$. Even though the gravitational torque remains between the mantle and inner core, the oscillation in this case is between the mantle and the whole of the TC (including the inner core). The larger moment of inertia of the whole of the TC, compared to the inner core alone, reduces the frequency of the MICG and lengthens its period. Using $\overline{\Gamma}=3\times 10^{20}$ N m as above, the period is now 5.9 years, approximately twice as long. If ¯$\overline{\cal F}_{i}$ is somewhere between these two extremes, with an amplitude such that the time it takes for $\Omega_{i}$ and $\Omega_{f}^{tc}$ to be brought back into co-rotation is of the order a few years, then the MICG mode is more complex, with each of $\Omega_{m}$, $\Omega_{i}$ and $\Omega_{f}^{tc}$ oscillating at different phases.
At periods of a few years, and assuming a radial magnetic field strength at the ICB of the order of 1 mT or larger, EM coupling is sufficiently strong to prevent any large differential rotation at the ICB (gubbins81). A rotation of the inner core will tend to entrain the fluid in contact with the ICB into co-rotation. This change in zonal flow near the ICB is communicated by the (strong) Coriolis force to the whole of the fluid column above and below the inner core, forcing it to rotate rigidly with it in accordance with the Proudman-Taylor constraint (jault08). Hence, on dynamical grounds, and for the expected strength of the magnetic field at the ICB, we expect that the MICG mode involves the whole of the TC, not just the inner core.
This simple presentation of the MICG mode is idealized in several ways. First, although ¯$\overline{\cal F}_{i}$ is most likely strong enough to lock the fluid cylinders inside the TC to the rotational motion of the inner core, some degree of differential rotation can occur near the outer edge of the TC (e.g. buffett96b; buffett98). Second, we have neglected the global elastic deformations that occur in response to the rotational displacement of the inner core out of its equilibrium longitudinal alignment. Taking these into account would lengthen the MICG period, though not by a large fraction. Third, if the inner core relaxes viscously over a relatively short time (approaching the MICG period), not only this can also lengthen the MICG period, it can lead to a significant attenuation of the mode. A fourth way this discussion of the MICG mode is idealized is that it assumes a complete decoupling of the fluid in the ROTC. Indeed, assessing the role of the ROTC on the MICG mode is the motivation for our study.
2.3 Torsional oscillation modes
The dynamical equation that governs the evolution of $\Omega_{f}$ is obtained by integrating the azimuthal component of the momentum equation over a cylindrical surface in the fluid core. For this specific geometry, the Coriolis acceleration term vanishes (e.g. taylor63; braginsky70), giving
$c_{f}\frac{\partial\Omega_{f}}{\partial t}=\frac{1}{\rho\mu}\frac{\partial}{\partial s}\left(c_{f}\{B_{s}\}\frac{b_{\phi}}{s}\right)-{\cal F}_{m}\left[\Omega_{f}-\Omega_{m}\right]-{\cal F}_{i}\left[\Omega_{f}-\Omega_{i}\right]+\nu\frac{\partial}{\partial s}\left(c_{f}\frac{\partial\Omega_{f}}{\partial s}\right)\,,$ (10)
@@ 2.3 Torsional oscillation modes
When both viscous and magnetic diffusion are neglected, inserting eq. (11) into eq. (10), we retrieve the characteristic equation for Alfvén waves subject to EM friction at the top and bottom of the cylinders, (e.g. braginsky70; buffett98)
$c_{f}\frac{\partial^{2}\Omega_{f}}{\partial t^{2}}=\frac{1}{\rho\mu}\frac{\partial}{\partial s}\left(c_{f}\{B_{s}^{2}\}\frac{\partial\Omega_{f}}{\partial s}\right)-{\cal F}_{m}\frac{\partial}{\partial t}\left[\Omega_{f}-\Omega_{m}\right]-{\cal F}_{i}\frac{\partial}{\partial t}\left[\Omega_{f}-\Omega_{i}\right]\,,$ (12)
where we have assumed $\{B_{s}\}^{2}\approx\{B_{s}^{2}\}$. The normal modes of oscillations of $\Omega_{f}$ that obey eq. (12) are the torsional oscillations (TO) modes (e.g. braginsky70). In their simplest form, when assuming ${\cal F}_{m}={\cal F}_{i}=0$, these modes describe the set of azimuthal oscillations of cylinders maintained by the magnetic tension between them. The terms that contain ${\cal F}_{m}$ and ${\cal F}_{i}$ capture the coupling of the fluid cylinders with the mantle (at the CMB) and the inner core (at the ICB). Keeping these terms not only provides a way by which TO modes can exchange angular momentum with the mantle and inner core, it also influences the structure of the modes.
In all previous studies on TO modes coupled with the mantle and inner core, eq. (12) was assumed for the dynamics of $\Omega_{f}$ (buffett96b; buffett98; mound03; mound05b; dumberry08c; dumberry10b). Here, we retain both magnetic and viscous diffusion and solve for both $\Omega_{f}$ and $b_{\phi}$ using the coupled eqs. (10) and (11). The motivation for retaining magnetic diffusion is grounded in the results of luo22a who showed that it reduces the quality factor of TO modes to values below 10. Although we expect $\nu\ll\eta$ in the fluid core, the motivation for retaining viscous diffusion in (10) is for numerical convenience, as it makes it easier to prescribe appropriate boundary conditions for $\Omega_{f}$. For all our results, we have used $\nu=\eta/100$, so as to ensure that magnetic diffusion dominates over viscous diffusion.
@@ 3 Numerical implementation
Equations (10) and (11) each require two boundary conditions, one each at $s_{o}$ and $s_{e}$. At $s=s_{o}$, we impose the regularity conditions $d\Omega_{f}/ds=0$ and $b_{\phi}=0$. At $s=s_{e}$, we set $d\Omega_{f}/ds=0$ which implies that there is no viscous torque between the fluid core and mantle and minimizes the influence of viscosity on our solutions. The condition on $b_{\phi}$ at $s_{e}$ is set by the magnetic field perturbation resulting from EM coupling (buffett98)
$b_{\phi}(s_{e})=-B_{s_{e}}\mu G_{m}s_{e}\left(\Omega_{f}(s_{e})-\Omega_{m}\right)\,,$ (14)
where $B_{s_{e}}$ is the rms amplitude of the radial magnetic field at the equator of the CMB.
@@
[…]
@@
We specify the location of the discrete $s$-grid points on two Chebyshev grids: one between $s_{o}$ and $s_{tc}$ and a second between $s_{tc}$ and $s_{e}$. The crowding of points near $s_{o}$, $s_{tc}$ and $s_{e}$ helps to resolve the thin viscous and magnetic boundary layers at $s_{o}$ and $s_{e}$ and to handle the rapidly changing $c_{f}$ in the region near the TC. We use 350 points inside the TC and 650 in the ROTC, for a total of $N_{s}$=1000 grid points, which proved to be sufficient to resolve the structure of the TO modes, and to ensure that angular momentum is conserved to within approximately 1 part in $10^{4}$.
@@
where ${\mathsf{A}}$ and ${\mathsf{B}}$ are $n\times n$ matrices. The resulting set of modes separate into a set of $i\omega$ with negative real part (decay) and a complex conjugate imaginary part (frequency); we only solve for the positive frequencies $\omega$. Note that ${\cal F}_{i}$ in eq. (5b) has a weak dependence on $\omega$ through the magnetic skin depth $\delta$. However, for a radial magnetic field strength at the ICB of 1 mT and above, the fluid cylinders inside the TC become quasi-locked to the inner core, and this frequency dependence only has a very small effect on the normal modes. To simplify, we use a nominal frequency of $62\pi/$ 6 yr-1 for calculating $\delta$, such that ${\cal F}_{i}$ is no longer a function of $\omega$, allowing us to write the system in the simple form of eq. (17). The same strategy has been employed in previous studies (e.g. buffett98; buffett05).
@@
$\{B_{s}\}=\overline{B}_{s}\sin\left(\frac{\pi(s-s_{o})}{s_{e}-s_{o}}\right)+\left(\frac{s-s_{o}}{s_{e}-s_{o}}\right)\frac{\overline{B}_{s}}{6}\,.$ (18)
The $\sin$ term produces a maximum field of $\overline{B}_{s}$ in the bulk of the region outside the TC, similar to the $B_{s}$ profile inverted by gillet10. The linear term ensures that the radial field at the equator does not vanish, and is set such that $B_{s_{e}}=\overline{B}_{s}/6$.
The numerical values that we adopt for all other parameters are given Table 1. The dipole field amplitude $B_{m}^{d}=0.319$ mT and the total rms strength of the radial field $\langle B_{r,m}\rangle=0.391$ mT at the CMB were obtained by taking the time-average of the gufm1 model of jackson00 over the time period 1950–1990 (the rms of the radial part of the dipole field alone is 0.226 mT).
Table 1: Parameters used in calculations
@@
(*) Based on a uniform fluid core density of $\rho=1.1\times 10^{4}$ kg m-3
@@ 4 Results
Let us first illustrate a case where the modes of axial oscillations between the mantle, inner core and fluid core include a distinct MICG mode. We set $\overline{\Gamma}=3\times 10^{20}$ N m, which gives an expected theoretical MICG frequency of $\omega_{micg}=1.065$ yr-1 (period of 5.902 years) based on eq. (9), and we set the mantle conductance at $G_{m}=10^{8}$ S. We set $\tau=100$ years to ensure that the MICG mode is only weakly damped by the viscous relaxation of the inner core. A key feature of the MICG mode is a weak involvement of the ROTC and this can be achieved by choosing a sufficiently weak $B_{s}$-field amplitude across the TC; we take $\overline{B}_{s}=0.1$ mT. We then compute the set of normal modes of oscillations for this setup, which we refer to as case $a$.
@@
Figure 1a shows the frequency and $Q$ of a subset of the resulting modes (light blue circles) that have frequencies in the vicinity of $\omega_{micg}$. For this choice of $\overline{B}_{s}$, the fundamental mode (first harmonic) of TO has a period of $\sim 180$ years; the sequence of modes that have a $Q$ between 3 and 4 on Figure 1a correspond to higher harmonics of TO. In addition to these, a distinct mode (identified by a purple outline) with a higher $Q$ of approximately 7, separated from the TO branch, is also present: this is the MICG mode. Its frequency matches well the expected theoretical prediction under the assumption that the whole of the TC co-rotates with the inner core. Figure 2a shows the profile of $\Omega_{f}$ for this MICG mode, at the phase when $\Omega_{i}$ is maximum in the prograde direction. Inside the TC, $\Omega_{f}$ is ocked with $\Omega_{i}$ except near the edge of the TC, as found in previous studies (buffett96a; buffett98) when the radial field strength at the ICB is above a few mT. The departure of $\Omega_{f}$ from $\Omega_{i}$ near the edge of the TC implies a slight reduction in the effective moment of inertia of the TC taking part in the MICG mode and explains why its frequency is slightly higher than its theoretical prediction (see Figure 1a). Also visible in Figure 2a are the Alfvén waves generated from the TC outward. With $\overline{B}_{s}=0.1$ mT, their wave speed is sufficiently slow that they get attenuated before they reach the equator of the CMB ($s=r_{f}$).
For comparison, we also show on Figure 1a the modes (black squares) that result when there is no gravitational coupling ($\overline{\Gamma}=0$). In this case, the MICG mode is absent from the spectrum of modes; only TO modes are present, and their frequency and $Q$ coincide almost identically with those found with $\overline{\Gamma}\neq 0$. These results show that when $B_{s}$ within the core is sufficiently weak, the MICG mode emerges as an independent mode of the system, along with a set of TO modes, and the latter are only weakly influenced by gravitational coupling.
Let us now increase $B_{s}$ while keeping $\overline{\Gamma}=3\times 10^{20}$ N m and $G_{m}=10^{8}$ S. We compute the modes for $\overline{B}_{s}=0.15$ mT (case $b$) and $\overline{B}_{s}=0.3$ mT (case $c$). The results are shown on Figure 1 (panels b and c, respectively), where we again show the comparison with the modes computed with $\overline{\Gamma}=0$. As $\overline{B}_{s}$ is increased, the MICG mode is no longer a distinct, independent mode of the system. Instead, it merges into the spectrum of TO modes, although the frequency and $Q$ of these TO modes is altered by gravitational coupling. To understand why, Figures 2b and 2c show the profiles of $\Omega_{f}$ for the modes close to the MICG frequency identified by a purple outline in panels $b$ and $c$ of Figure 1. With the larger $B_{s}$-field, Alfvén waves generated at the TC travel faster and can propagate further before they get attenuated. While for case $b$ Alfvén waves are just barely able to make it through the entire width of the ROTC, they easily reach $s=r_{f}$ in case $c$. Consequently, the structure of Alfvén waves within the whole of the ROTC must be compatible with their end-condition at $s=r_{f}$ and they thus have a feedback on the oscillating motion of the TC. In other words, the oscillating TC is now taking part in TO modes, although these are modified by gravitational coupling.
@@
Figure 1: Frequency and $Q$ of a subset of the normal modes obtained with (a) $\overline{B}_{s}=0.1$, (b) $\overline{B}_{s}=0.15$, and (c) $\overline{B}_{s}=0.3$ and different choices of coupling parameters (Pale blue circles: $\overline{\Gamma}=3\times 10^{20}$ N m, $G_{m}=10^{8}$ S; black squares: $\overline{\Gamma}=0$ N m, $G_{m}=10^{8}$; red diamonds: $\overline{\Gamma}=3\times 10^{20}$ N m, $G_{m}=10^{9}$ S). In each panel, the predicted $\omega_{micg}$ from Eq. (9) is indicated by the vertical orange line. The angular velocity structure of the modes identified by a purple outline is shown in Figure 2.
The attenuation of Alfvén waves can be achieved by reducing their speed, as illustrated by case $a$, or by increasing their attenuation rate from EM coupling at the CMB. To illustrate this, we carried out a set of experiments where we have kept $\overline{\Gamma}=3\times 10^{20}$ N m, the same three values of $\overline{B}_{s}$ as for cases $a$, $b$ and $c$ above, but used an increased mantle conductance of $G_{m}=10^{9}$ S. The modes that result from these experiments are shown on Figure 1 as red diamonds. The $Q$ of all modes is reduced because of the enhanced attenuation caused by the stronger EM coupling at the CMB. In each case though, a distinct MICG mode remains present in the mode spectrum, even with $\overline{B}_{s}=0.3$ mT.
The above analysis allows us to refine the criteria that determines whether the MICG remains an independent mode of the system. The condition is whether Alfvén waves launched by the oscillating TC get attenuated before they can reach $s=r_{f}$. When Alfvén waves are attenuated, they do not have a feedback on the motion of the TC, and their oscillating nature imply a weak net angular momentum change of the ROTC. In this sense, the ROTC plays a passive role, and the MICG mode, for which the angular momentum exchange remains dominantly between the mantle and the TC via gravitational coupling, is a possible independent mode of oscillation. In contrast, if Alfvén waves launched at the TC can travel through the entire width of the ROTC, the motion of the TC must be compatible with the allowed structure of TO modes.

### v2

@@ 2.3 Torsional oscillation modes
When both viscous and magnetic diffusion are neglected, inserting eq. (11) into eq. (10), we retrieve the characteristic equation for Alfvén waves subject to EM friction at the top and bottom of the cylinders, [4, 8, e.g.]
[…]
where we have assumed $\{B_{s}\}^{2}\approx\{B_{s}^{2}\}$. The normal modes of oscillations of $\Omega_{f}$ that obey eq. (12) are the torsional oscillations (TO) modes [4, e.g.]. In their simplest form, when assuming ${\cal F}_{m}={\cal F}_{i}=0$, these modes describe the set of azimuthal oscillations of cylinders maintained by the magnetic tension between them. The terms that contain ${\cal F}_{m}$ and ${\cal F}_{i}$ capture the coupling of the fluid cylinders with the mantle (at the CMB) and the inner core (at the ICB). Keeping these terms affects the structure of the modes and provides a way by which TO modes can exchange angular momentum with the mantle and inner core.
In most previous studies on TO modes coupled with the mantle and inner core, eq. (12) was assumed for the dynamics of $\Omega_{f}$ [6, 8, 37, 38, 20, 21]. The studies of [5] and [40] included the effects of viscous diffusion both within the volume of the fluid core and at boundaries. Here, we retain both magnetic and viscous diffusion and solve for both $\Omega_{f}$ and $b_{\phi}$ using the coupled eqs. (10) and (11). Although we expect $\nu\ll\eta$ in the fluid core, the motivation for retaining viscous diffusion in (10) is for numerical convenience, as it makes it easier to prescribe appropriate boundary conditions for $\Omega_{f}$. For all our results, we have used $\nu=\eta/100$, so as to ensure that magnetic diffusion dominates over viscous diffusion. Although it would be preferable to solve for $\Omega_{f}$ and $b_{\phi}$ as a function of both $s$ and $z$ [35, e.g.], our simple one-dimensional (1D) model captures the leading order dynamics of Alfvén waves within the core.
As we shall show, whether the MICG mode is an independent mode of the core-mantle system depends on the travel time versus diffusion time of Alfvén waves in the ROTC. Alfvén waves travel at velocity $V_{A}=\{B_{s}\}/\sqrt{\rho\mu}$, and the time they take to travel between the TC and the equator of the CMB ($s=r_{f}$) is $\tau_{A}=(r_{f}-r_{i})/V_{A}$. Their typical timescale of diffusion from ohmic dissipation within the core is $\tau_{d}=1/(k_{s}^{2}\eta)$, where $k_{s}$ is the wavenumber in the $s$-direction. Alfvén waves obey the dispersion relation $\omega=k_{s}V_{A}$, so TO modes that have frequencies close to MICG mode should have a typical wavenumber $k_{s}=\omega_{micg}/V_{A}$, giving $\tau_{d}=V_{A}^{2}/\eta\omega_{micg}^{2}$. Alfvén waves are also attenuated by EM coupling at the CMB. A simple estimate of this EM attenuation timescale can be obtained from eq. (10) and is $\tau_{em}=c_{f}/{\cal F}_{m}$. This quantity depends on $s$ but a rough bulk estimate is given by $\tau_{em}=\rho r_{c}/(G_{m}\langle B_{r,m}\rangle^{2})$. The travel time versus diffusion time of Alfvén waves is then characterized by two Lundquist numbers,
$Lu=\frac{\tau_{d}}{\tau_{A}}\,,\qquad Lu_{em}=\frac{\tau_{em}}{\tau_{A}}\,.$ (13)
If either $Lu<1$ or $Lu_{em}<1$, Alfvén waves are attenuated before they travel the full width of the ROTC.
Another useful diagnostic quantity is the Lundquist number based on the travel versus diffusion time of Alfvén waves in the conducting mantle,
$Lu_{m}=\sqrt{\frac{\mu}{\rho}}\,G_{m}\langle B_{r,m}\rangle\,.$ (14)
In particular, $Lu_{m}$ determines the degree of reflection of Alfvén waves when they reach the equator of the CMB; when $Lu_{m}\approx 1$, Alfvén waves are fully absorbed [44, 24]. This quantity is denoted as $Q$ in both [44] and [24], but here we reserve $Q$ to denote the quality factor of modes.
@@ 3 Numerical implementation
Eqs. (10) and (11) each require boundary conditions at $s_{o}$ and $s_{e}$. At $s=s_{o}$, we impose the regularity conditions $d\Omega_{f}/ds=0$ and $b_{\phi}=0$. At $s=s_{e}$, the condition on $b_{\phi}$ is set by the magnetic field perturbation resulting from EM coupling [8, e.g. eq. 12 of]
$b_{\phi}(s_{e})=-\langle B_{r,m}\rangle\mu G_{m}s_{e}\left(\Omega_{f}(s_{e})-\Omega_{m}\right)\,$, (16)
and we impose a free-slip condition, $d\Omega_{f}/ds=0$. The latter implies that there is no viscous torque between the fluid core and mantle and minimizes the influence of viscosity on our solutions. Since we use $\nu=\eta/100$, the free-slip condition is met by a flow adjustment in a thin viscous boundary layer near $s=s_{e}$.
@@
$\hat{\Gamma}_{cmb}$ $=\int_{s_{o}}^{s_{e}}{\cal F}_{m}(s)\left[\Omega_{f}-\Omega_{m}\right]ds$
$\hskip 14.22636pt+4\pi s_{e}^{3}\sqrt{1-s_{e}^{2}}\,G_{m}\langle B_{r,m}\rangle^{2}\left(\Omega_{f}(s_{e})-\Omega_{m}\right)\,$, (18a)
$\hat{\Gamma}_{icb}$ $=\int_{s_{o}}^{s_{tc}}{\cal F}_{i}(s)\left[\Omega_{f}-\Omega_{i}\right]ds\,.$ (18b)
@@
We specify the location of the discrete $s$-grid points on two Chebyshev grids: one between $s_{o}$ and $s_{tc}$ and a second between $s_{tc}$ and $s_{e}$. The crowding of points near $s_{o}$, $s_{tc}$ and $s_{e}$ helps to resolve the thin viscous and magnetic boundary layers at $s_{o}$ and $s_{e}$ and to handle the rapidly changing $c_{f}$ in the region near the TC. We use 350 points inside the TC and 650 in the ROTC, for a total of $N_{s}$=1000 grid points, which proved to be sufficient to resolve the structure of the relevant TO modes, and to ensure that angular momentum is conserved to within approximately 1 part in $10^{4}$.
@@
where $\mathsf{A}$} and $\mathsf{B}$} are $n\times n$ matrices. The solutions are the possible free modes of the system. The imaginary part of $\omega$ is negative, capturing mode decay. The real part of $\omega$ is either positive or negative (with the same magnitude) and represents the frequency of oscillation; we only solve for the positive frequencies. Note that ${\cal F}_{i}$ in eq. (5b) has a weak dependence on $\omega$ through the magnetic skin depth $\delta$. However, for a radial magnetic field strength at the ICB $\geq 1$ mT, the fluid cylinders inside the TC become quasi-locked to the inner core, and this frequency dependence only has a very small effect on the normal modes. To simplify, we use a nominal frequency of $2\pi/6$ yr-1 for calculating $\delta$, such that ${\cal F}_{i}$ is no longer a function of $\omega$, allowing us to write the system in the simple form of eq. (19). The same strategy has been employed in previous studies [8, 10, e.g.].
@@
$\{B_{s}\}=\overline{B}_{s}\sin\left(\frac{\pi s}{r_{f}}\right)+\left(\frac{s}{r_{f}}\right)B_{s_{e}}\,$, (20)
where $\overline{B}_{s}$ is an amplitude factor and $B_{s_{e}}$ is the amplitude of the radial magnetic field at the equator of the CMB. The sine term produces a maximum $B_{s}$-field in the middle of the core, similar to the profile inverted by [25]. The linear term ensures that the radial field at the equator does not vanish. For our numerical experiments in the weak-field regime, $\overline{B}_{s}$ is varied between 0.1-0.3 mT, and we set $B_{s_{e}}=\overline{B}_{s}/6$. In the strong-field regime, we set $\overline{B}_{s}=3.3$ mT and $B_{s_{e}}=\langle B_{r,m}\rangle$.
The numerical values that we adopt for all other parameters are given in Table 1. The dipole field amplitude $B_{m}^{d}=0.319$ mT and the total rms strength of the radial field $\langle B_{r,m}\rangle=0.391$ mT at the CMB were obtained by taking the time-average of the gufm1 model of [31] over the time period 1950–1990 (the rms of the radial part of the dipole field alone is 0.226 mT). We note that, for our experiments in the weak-field regime, there is an inconsistency between our chosen ICB and CMB field strengths and the lower ${B}_{s}$-field in the range of 0.1-0.3 mT. These low values were chosen because they are similar to those used in previous studies of the MICG mode [37, 39, e.g.] and they help to illustrate the role that Alfvén waves have on this mode.
We show results for different choices of mantle conductance: $G_{m}=10^{8}$ S, $2\times 10^{8}$ S, $4\times 10^{8}$ S, and $6\times 10^{8}$ S. The corresponding $Lu_{m}$ values, in increasing order of $G_{m}$, are: 0.34, 0.68, 1.36, 2.05. These remain sufficiently low that our 1D model captures correctly the reflection properties of Alfvén waves as they reach the equator of the CMB [44, 24].
Table 1: Parameters used in calculations.
@@
(*) Based on a uniform fluid core density of $\rho=1.1\times 10^{4}$ kg m-3.
@@ 4 Results
Let us first illustrate a case where the modes of axial oscillations between the mantle, inner core and fluid core include a distinct MICG mode. We set $\overline{\Gamma}=3\times 10^{20}$ N m, which gives an expected MICG frequency of $\omega_{micg}=1.065$ yr-1 (period of 5.902 years) based on eq. (9), and we set the mantle conductance at $G_{m}=10^{8}$ S. We set $\tau=100$ years to ensure that the MICG mode is only weakly damped by the viscous relaxation of the inner core. We choose $\overline{B}_{s}=0.1$ mT, which gives a typical travel time for Alfvén waves across the ROTC of $\tau_{A}=84$ yr. We compute the set of normal modes of oscillations for this setup, which we refer to as case $a$.
@@
Figure 1a shows the frequency and quality factor $Q$ of a subset of the resulting modes (light blue circles) that have frequencies in the vicinity of $\omega_{micg}$. For this choice of $\overline{B}_{s}$, the fundamental mode (first harmonic) of TO has a period of $\sim 180$ years; the sequence of modes that have a $Q$ between 3 and 4 on Figure 1a correspond to higher harmonics of TO. In addition to these, a distinct mode (identified by a purple outline) with a higher $Q$ of approximately 7, separated from the TO branch, is also present: this is the MICG mode. Its frequency matches well the expected theoretical prediction under the assumption that the whole of the TC co-rotates with the inner core. Figure 2a shows the profile of $\Omega_{f}$ for this MICG mode, at the phase when $\Omega_{i}$ is maximum in the prograde direction. Inside the TC, $\Omega_{f}$ is locked with $\Omega_{i}$ except near the edge of the TC, as found in previous studies [7, 8]. The rotation of the mantle, $\Omega_{m}$, is opposite that of the inner core, and its amplitude matches that of the MICG mode under the assumption that the whole of the TC oscillates as a rigid body, $\Omega_{m}=-(C_{tc}/C_{m})\Omega_{i}$. (Note that $\Omega_{m}$ is increased by a factor 100 on Figure 2 so that it reaches a similar scale as $\Omega_{i}$ and $\Omega_{f}$). The departure of $\Omega_{f}$ from $\Omega_{i}$ near the edge of the TC implies a slight reduction in the effective moment of inertia of the TC taking part in the MICG mode and explains why its frequency is slightly higher than its theoretical prediction (see Figure 1a). Also visible in Figure 2a are the Alfvén waves generated from the TC outward. With $\overline{B}_{s}=0.1$ mT, and at the frequency of the MICG mode, their wavelength is sufficiently small that their typical time of attenuation is $\tau_{d}=25$ yr. This is shorter than $\tau_{A}$, so $Lu<1$, and Alfvén waves get attenuated before they reach the equator of the CMB ($s=r_{f}$). These waves carry angular momentum but their short wavelengths imply a small net angular momentum when integrated over the ROTC, and the latter plays a negligible role in the MICG angular momentum budget.
@@
Figure 1: Frequency and $Q$ of a subset of the normal modes obtained with $\overline{\Gamma}=3\times 10^{20}$ N m, (a) $\overline{B}_{s}=0.1$, (b) $\overline{B}_{s}=0.15$, and (c) $\overline{B}_{s}=0.3$, and different choices of mantle conductance $G_{m}$ (see legend). Black squares show the TO modes in the absence of gravitational coupling: $\overline{\Gamma}=0$. In each panel, the predicted $\omega_{micg}$ from Eq. (9) is indicated by the pink vertical line. The angular velocity structure of the modes identified by a purple outline is shown in Figure 2.
For comparison, we also show on Figure 1a the modes (black squares) that result when there is no gravitational coupling ($\overline{\Gamma}=0$). In this case, the MICG mode is absent from the spectrum of modes; only TO modes are present, and their frequencies and $Q$ coincide almost identically with those found with $\overline{\Gamma}\neq 0$. These results show that when $B_{s}$ within the core is sufficiently weak, the MICG mode emerges as an independent mode of the system, along with a set of TO modes, and the latter are only weakly influenced by gravitational coupling.
For a mantle conductance of $G_{m}=10^{8}$, the typical EM attenuation time of Alfvén waves is $\tau_{em}=119$ yr. We also show on Figure 1a the set of modes for three other choices of $G_{m}$: $2\times 10^{8}$ S ($\tau_{em}=60$ yr), $4\times 10^{8}$ S ($\tau_{em}=30$ yr) and $6\times 10^{8}$ S ($\tau_{em}=20$ yr). The $Q$ of all modes is reduced because of the enhanced attenuation from EM coupling at the CMB. In each case, the MICG mode is present as an independent mode because $Lu<1$, while $Lu_{em}$ is also <1<1 for the 3 cases with higher mantle conductance.
Let us now increase $B_{s}$ while keeping $\overline{\Gamma}=3\times 10^{20}$ N m and restoring $G_{m}=10^{8}$ S. We compute the modes for $\overline{B}_{s}=0.15$ mT (case $b$) and $\overline{B}_{s}=0.3$ mT (case $c$). The results are shown on Figure 1 (panels b and c, respectively), where we again show the comparison with the modes computed with $\overline{\Gamma}=0$. Figures 2b and 2c show the profiles of $\Omega_{f}$ for the modes close to the MICG frequency identified by a purple outline in panels $b$ and $c$ of Figure 1. With a larger $B_{s}$-field, Alfvén waves generated at the TC travel faster, their wavelengths at frequencies close to $\omega_{micg}$ is longer (increasing their ohmic diffusion time $\tau_{d}$), so they propagate further before they get attenuated. The travel and ohmic diffusion times for case $b$ are $\tau_{A}=56$ yr, $\tau_{d}=58$ yr and those for case $c$ are $\tau_{A}=28$ yr, $\tau_{d}=235$ yr. In each case, $Lu>1$ (though barely for case $b$) and (with $\tau_{em}=119$ yr) $Lu_{em}>1$; Alfvén waves can travel the width of ROTC before they get attenuated. Consequently, the structure of Alfvén waves within the whole of the ROTC must be compatible with their end-condition at $s=r_{f}$ and they thus have a feedback on the oscillating motion of the TC. In each case, the MICG mode is no longer a distinct, independent mode of the system. Instead, it merges into the spectrum of TO modes.
Figures 1b and 1c also show the resulting modes when the mantle conductance is increased to $2\times 10^{8}$ S ($\tau_{em}=60$ yr), $4\times 10^{8}$ S ($\tau_{em}=30$ yr) and $6\times 10^{8}$ S ($\tau_{em}=20$ yr). By progressively decreasing $\tau_{em}$, a point is reached when Alfvén waves get attenuated before they can travel the width of the ROTC (that is, $Lu_{em}$ becomes smaller than 1), and a distinct MICG mode re-emerges in the spectrum of modes. With $\overline{B}_{s}=0.15$ mT ($\tau_{A}=56$ yr), the MICG is almost fully out of the TO branch for $G_{m}=2\times 10^{8}$ S ($Lu_{em}\sim 1$), and clearly separated for $G_{m}>4\times 10^{8}$ S . With $\overline{B}_{s}=0.3$ mT ($\tau_{A}=28$ yr), a larger $G_{m}=6\times 10^{8}$ S is required for $Lu_{em}<1$ and for the MICG mode to split from the TO branch.